Determine the eight term (simplified) in the expansion of (4a+7b)^15.
Thank you
Kristen

Answer by stanbon(75887)

Determine the eight term (simplified) in the expansion of (4a+7b)^15.
8th term = 15C8(4a)^8(7b)^7 = 6435*65636*a^8*823543b^7=(3.4784x10^14)a^8b^7
-------
Cheers,
Stan H.